OpenStudy (anonymous):

Evaluate the Integral: (x^2tanx)/(1 + cos^4x) dx [-pi/4, pi/4]

OpenStudy (anonymous):

is 2tanx the exponent?

OpenStudy (anonymous):

no 2 is

OpenStudy (anonymous):

For everyone else:\[\large\int_{-\frac{\pi}{4}}^{\frac{\pi}{4}}\frac{x^2\tan x}{1+\cos^4x}dx\]

OpenStudy (anonymous):

thanks that's exactly my question

OpenStudy (anonymous):

Is it an odd or even function?

OpenStudy (anonymous):

When a function is odd, \[ \int_{-a}^a f(x)dx = 0 \]

OpenStudy (anonymous):

yeah your right... that would make sense, it is an odd function... thank u :)